OpenStudy (anonymous):

y varies inversely with x, and x=150 when y=30. Find the inverse variation equation and use it to find the value of x when y=90 a. 1/5 b. 9/5 c. 5 d. 5/9 PLEASE HELP ME!

OpenStudy (anonymous):

y=k/x to find k: 30=k/150 k=30*50=1500 so equation is: y=1500/x now finf value of x, when y =90 90=1500/x x=1500/90 = 150/9

OpenStudy (anonymous):

so what that equal too? 5/9???

OpenStudy (anonymous):

150/9=50/3

OpenStudy (anonymous):

than what?

OpenStudy (anonymous):

that's it

OpenStudy (anonymous):

but it doesn't match to any of the answers though

OpenStudy (anonymous):

ups, i made a mistake, :) y=k/x to find k: 30=k/150 k=30*150=4500 so equation is: y=4500/x now finf value of x, when y =90 90=4500/x x=4500/90 = 450/9=50

OpenStudy (anonymous):

so then is it d??

OpenStudy (anonymous):

no, it's x=50, but it doesn't mach any answer. Are you sure that x=150 when y=30 and no x=15 when y=30

OpenStudy (anonymous):

yeah i'm sure...

OpenStudy (anonymous):

so x=50 is the answer

OpenStudy (anonymous):

so c???

OpenStudy (anonymous):

the most probable is c, becouse i think there is a mistake in the problem